General Electric has just issued a callable​ (at par)​ 10-year, 6.5% coupon bond with annual coupon payments. The bond can be called at par in one year or anytime thereafter on a coupon payment date. It has a price of $101.77.

a. What is the​ bond's yield to​ maturity?

b. What is its yield to​ call?

c. What is its yield to​ worst?
Answer #1

a]

YTM is calculated using RATE function in Excel with these inputs :

nper = 10 (10 years to maturity with 1 annual coupon payment each year)

pmt = 100 * 6.5% (annual coupon payment = face value * annual coupon rate. This is a positive figure as it is an inflow to the bondholder)

pv = -101.77 (current bond price. This is a negative figure as it is an outflow to the buyer of the bond)

fv = 100 (face value of the bond receivable on maturity. This is a positive figure as it is an inflow to the bondholder)

The RATE is calculated to be 6.26%. This is the YTM

х v f =RATE(10,100*6.5%,-101.7,100) B C D A 6.26%|

b]

YTC is calculated using RATE function in Excel with these inputs :

nper = 1 (1 years to first call date with 1 annual coupon payment each year)

pmt = 100 * 6.5% (annual coupon payment = face value * annual coupon rate. This is a positive figure as it is an inflow to the bondholder)

pv = -101.77 (current bond price. This is a negative figure as it is an outflow to the buyer of the bond)

fv = 100 (call price of the bond receivable on call date. This is a positive figure as it is an inflow to the bondholder)

The RATE is calculated to be 4.65%. This is the YTC.

A1 X fc C =RATE(1,100*6.5%,-101.77,100) D E F G A B 4.65%!

c]

Yield to worst is the lower of YTM and YTC

Yield to worst is 4.65%